OpenStudy (anonymous):

The eighteenth term of an arithmetic sequence is 243, and the ninetieth term is 516.6. What is the 125th term?

OpenStudy (dumbcow):

\[a _{n} = a _{1} + d(n-1)\] \[d = 516.6 - 243 = 273.6\] \[a _{18} = 243 = a _{1} + 273.6(18-1)\] \[\rightarrow a _{1} = -4408.2\] \[a _{125} = -4408.2 + (273.6)(124) = 29518.2\]